# Performance-Portable Binary Neutron Star Mergers with AthenaK

### v2

@@ 1 Introduction
AthenaK 111see https://github.com/IAS-Astrophysics/athenak is a complete rewrite of the Athena++ code using Kokkos. This paper is the third in a series describing its features and capabilities, with Stone & Mullen (2024) focusing on its architecture and applications to astrophysical fluids, Zhu et al. (2024) discussing the dynamical spacetime solver and infrastructure for binary black hole (BBH) problems, and this paper discussing an extension which couples GRMHD to dynamical spacetimes. In Section 2, we briefly discuss the GRMHD equations. Section 3 provides an overview of our numerical methods, with particular attention given to our atmosphere treatment and the implementation of a first-order flux correction (FOFC) scheme. We demonstrate the code’s accuracy and robustness in Section 4, including a convergence test with the first published BNS mergers performed on GPUs. Lastly, Section 5 describes the results of our performance and scaling tests, demonstrating AthenaK’s suitability for exascale machines. Lastly, we summarize our work in Section 6 and highlight some potential new NR applications for AthenaK.
@@ 3 Numerical Methods
In this section, we provide an overview of the numerical methods we use for the Valencia GRMHD solver. For an explanation of the refinement structure and methods shared with other modules in AthenaK, we refer the reader to Stone & Mullen (2024).
@@ 3.3 Atmosphere Treatment
3.4 Primitive Variable Recovery
4 Tests
In this section, we describe a number of tests to validate AthenaK.
4.1 Magnetized Shock Tubes
@@ 3 $L$ $5/3$ $1$ $10^{3}$ 00 00 00 $10$ $7$ $7$
Figure 1: The rest-mass density $\rho$, pressure $P$, $x$-velocity $u^{x}=Wv^{x}$, and $y$-component of the magnetic field $B^{y}$ for Balsara Test #1 at $t=0.5$
Figure 2: The same as Figure 1, but for Balsara Test #2.
Figure 3: The same as Figure 1, but for Balsara Test #3.
Figure 4: The same as Figure 1, but for Balsara Test #4.
Figure 5: The same as Figure 1, but for Balsara Test #5.
Our computational domain contains $1600$ cells. We use the LLF Riemann solver with PPM4 reconstruction. Because the fluid solver remains stable during this test and never approaches atmosphere, we do not enable the FOFC scheme. We evolve these systems to $t=0.5$ and compare them to their corresponding exact solutions (Giacomazzo & Rezzolla, 2006) in Figures 15. For brevity, we only show the $x$-component of velocity and the $y$-component of the magnetic field. The results are in good agreement with the exact solution.
@@ 4.2 Magnetized Cylindrical Blast Wave
For this test, we set $\rho_{\textrm{amb}}=10^{-4}$, $P_{\textrm{amb}}=3\times 10^{-5}$, $\rho_{\textrm{in}}=10^{-2}$, $P=\textrm{in}=1$, $r_{\textrm{in}}=0.8$, and $r_{\textrm{out}}=1.0$. We consider three values for the magnetic field: $B_{x}=0.01$, $B_{x}=0.1$, and $B_{x}=1.0$. The EOS is an ideal gas with $\Gamma=4/3$. As noted in Beckwith & Stone (2011), this test is particularly difficult for large $B_{x}$ because the combination of relatively high Lorentz factors and high magnetizations heavily often causes the conserved-to-primitive inversion to return poor-quality solutions (if it finds them at all). Common solutions to this issue involve altering the test (e.g., increasing the ambient pressure as in Leismann et al. (2005); Del Zanna et al. (2007); Beckwith & Stone (2011); Deppe et al. (2023)) or employing schemes which violate energy conservation (Komissarov, 1999; Mignone & Bodo, 2006; Phillips & Komissarov, 2024; Komissarov & Phillips, 2024). Consequently, the original setup as in Komissarov (1999) tests both the robustness of our primitive recovery routine and the ability of the FOFC and FOFC+DMP schemes to add enough dissipation to limit unphysical behavior.
@@
We show the results for the $B_{x}=0.01$ case in Figure 7, the $B_{x}=0.1$ case in Figure 8, and $B_{x}=1.0$ in Figure 9. Both the FOFC and FOFC+DMP schemes are in good agreement for the weakly and moderately magnetized cases. The strongly magnetized case exhibits significant noise for both solutions, with FOFC+DMP offering somewhat better behavior near the center. Nevertheless, the overall solution agrees with Komissarov (1999), and our results are consistent with what was found for the same test using the HARM-like solver (Stone & Mullen, 2024).
@@ 4.3 Magnetic Loop Advection
Figure 10: (Top) The $B^{x}$ component of the magnetic field during the loop advection test. (Bottom) The violation of the divergence-free constraint. In both plots, the left half shows the initial data, and the right half shows the evolved data after a single period ($t=2.4$).
This test probes the correctness and the diffusiveness of the constrained-transport scheme. As shown in Figure 10, AthenaK advects the loop at the correct speed. Some numerical diffusion is apparent, but the shape is correct with no obvious artifacts. Maximum errors in the divergence-free condition are around $7\times 10^{-17}$ and primarily concentrated around the edge of the loop. After a single period, the maximum error has amplified to approximately $3\times 10^{-15}$, which is consistent with growth from floating-point errors.
@@ 4.4 Oscillating Neutron Star in the Cowling Approximation
The computational domain spans $[0~{}\textrm{M}_{\odot},102.4~{}\textrm{M}_{\odot}]$ in all dimensions with vacuum boundary conditions and a reflection symmetry along the $x$, $y$, and $z$ axes. The base refinement level has $128$ cells in each direction, and we use SMR to refine all mesh blocks inside $51.2~{}\textrm{M}_{\odot}$ and then again at $25.6~{}\textrm{M}_{\odot}$ to give the innermost refinement region a resolution of $0.2~{}\textrm{M}_{\odot}$. We set the Courant-Friedrich-Lewy (CFL) factor to $\textrm{CFL}=0.4$.
@@ 4.6 Magnetized Accretion Disk
Figure 13: Diagnostic fluxes for the $64^{3}$ case. For comparison, some of the $96^{3}$ EHT data is also plotted. From top to bottom, we show the rates for mass accretion, normalized energy accretion, normalized angular momentum accretion, and normalized magnetic flux.
Figure 14: The same as Figure 13 but for the $64^{3}$ case, which is compared to the $128^{3}$ EHT data.
Figure 15: The same as Figure 13 but for the $128^{3}$ case, which is compared to the $192^{3}$ EHT data.
@@
We show these diagnostic quantities in Figures 13, 14, and 15 for the $64^{3}$, $96^{3}$, and $128^{3}$ cases, respectively. For comparison, the plots also contain a subset of the data from Porth et al. (2019) (hereafter referred to as the EHT data), namely Athena++, BHAC, Cosmos++ (Anninos et al., 2005; Fragile et al., 2012, 2014), ECHO, HARM-Noble (usually called HARM3d) and iharm3D (Gammie et al., 2003; Noble et al., 2006, 2009), IllinoisGRMHD, and KORAL (Sadowski et al., 2013, 2014). Somewhat confusingly, we compare the $64^{3}$ AthenaK runs to the $96^{3}$ EHT data, the $96^{3}$ runs to the $128^{3}$ EHT data, and the $128^{3}$ runs to the $192^{3}$ EHT data. This is because the AthenaK grids do not have a direct correspondence with any setup in the EHT data; all AthenaK setups are Cartesian, while most of the EHT data setups are on logarithmic spherical grids. Nevertheless, we see good agreement between AthenaK and the other codes present at all resolutions, particularly for the angular momentum flux and the mass accretion rates.
@@ 4.7 Binary Neutron Star Merger
We use the SR and HR resolution runs to establish convergence of the waveform phase. To eliminate phase offsets caused by non-converging noise prior to the arrival of the so-called “junk” radiation at the extraction radius, we align the waveforms using the procedure in Boyle et al. (2008) between $t-r_{\ast}=0$ and the merger time of the LR waveform. We show the results in the bottom plot in Figure 21. Throughout the inspiral and merger, we maintain approximate second-order convergence as expected by our numerical methods. The results do not show clean convergence in the post-merger phase even before collapse, suggesting that we need at least one more higher resolution run.
@@
We next consider the reduced binding energy, $E_{b}=(M_{\mathrm{ADM}}-E_{\mathrm{GW}}-M)/(M\nu)$, as a function of the reduced angular momentum, $j=(J_{\mathrm{ADM}}-J_{\mathrm{GW}})/(M^{2}\nu)$, where for our initial data $M_{\mathrm{ADM}}=2.9984~{}\textrm{M}_{\odot}$, $J_{\mathrm{ADM}}=8.83542~{}\textrm{M}_{\odot}^{2}$ in natural units, and the symmetric mass ratio is $\nu=q/(1+q)^{2}=1/4$. This is a gauge-invariant measure which shows the evolution of the binary as it loses both energy and angular momentum due to gravitational radiation (Damour et al., 2012; Bernuzzi et al., 2012). The bottom plot in Figure 22 shows that our runs are all in excellent agreement prior to merger. The lines do diverge somewhat in the post-merger phase, but the differences in the AthenaK runs decrease with resolution, indicating some measure of convergence.
@@ 5.2 Strong Scaling
We show the strong-scaling test results in Figure 25. Relative to a single node (${\sim}236^{3}$ zones per GPU), AthenaK on Frontier maintains $\geq 80\%$ efficiency down to ${\sim}96^{3}$ zones per GPU for the TOV test. The $64^{3}$ block test is roughly twice as fast as the $32^{3}$ test, suggesting that $32^{3}$ mesh blocks do not saturate the GPUs properly. Despite the performance difference, scaling results are comparable between the two setups at smaller node counts. However, there is some evidence suggesting that large mesh blocks scale more poorly at large node counts. Though the limited tests shown here are hardly conclusive, this seems reasonable; larger mesh blocks improve GPU vectorization and reduce the total number of communication calls, but the cost of launching a kernel does not substantially change, meaning that kernel overhead will become a bottleneck sooner than in the case of small mesh blocks.
@@ 5.3 Discussion on Performance
In addition to the scaling results above, we performed a small number of tests comparing CPU and GPU performance on NERSC Perlmutter, which consists of CPU nodes (each with two AMD EPYC 7763 CPUs) and GPU nodes (each with a single CPU and four Nvidia A100 GPUs). Our setup is similar to Sec. 4.5 but with all output disabled and limited to the first 20 time steps. Running with a single CPU (64 cores/128 threads) achieves $7.36\times 10^{5}$ zone-cycles/s, and a single Nvidia A100 GPU manages to achieve $8.72\times 10^{6}$ zone-cycles/s, a factor of ${\sim}12$ improvement. To provide a baseline comparison, we also perform this test with GR-Athena++ using as similar a setup as possible. For a cell-centered spacetime, GR-Athena++ achieves $9.64\times 10^{5}$ zone-cycles/s, which is somewhat faster but comparable to AthenaK’s CPU performance.
We also ran a version of this test on a consumer-grade laptop with an Nvidia RTX 3070 laptop GPU. Due to the reduced amount of memory, we reduced the resolution by half, switched to $16^{3}$ mesh blocks, and imposed a reflection symmetry across the z-axis, but the physics options are the same. The laptop GPU achieves $1.03\times 10^{6}$ zone-cycles/s, which is $40\%$ faster than a single CPU on Perlmutter. The same calculation using the laptop’s CPU (an AMD Ryzen 9 5900X) with 16 OpenMP threads only achieved $1.25\times 10^{5}$ zone-cycles/s. We also found it possible to run a very coarse (${\sim}900~{}\mathrm{m}$) BNS simulation through merger on this same device in a matter of a few hours. Though research-grade calculations still require larger machines, the order-of-magnitude speedup provided by GPUs makes it possible to accelerate testing and debugging considerably.
@@ 6 Conclusion
The authors thank the Penn State Numerical Relativity Group for helpful discussions and important feedback. This research makes use of data from the EHT Code Comparison Project, and JF thanks Oliver Porth for assistance with accessing the data. SB and BD acknowledge support by the EU Horizon under ERC Consolidator Grant, no. InspiReM-101043372. This research was supported by funding from the U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Number(s) DE-SC0021177 and DE-SC0024388, by NASA under award No. 80NSSC21K1720, and by the National Science Foundation under Grants No. PHY-2011725, PHY-2020275, PHY-2116686, and AST-2108467. Part of this work was completed during a sabbatical visit to the Institute for Gravitation and the Cosmos at the Pennsylvania State University by SB. Simulations were performed on OLCF’s Frontier, NERSC’s Perlmutter, and on the Pennsylvania State University’s Institute for Computational and Data Sciences’s Roar Collab supercomputer. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.

### v3

@@ 1 Introduction
AthenaK 111see https://github.com/IAS-Astrophysics/athenak is a complete rewrite of the Athena++ code using Kokkos. This paper is the third in a series describing its features and capabilities, with Stone et al. (2024) focusing on its architecture and applications to astrophysical fluids, Zhu et al. (2024) discussing the dynamical spacetime solver and infrastructure for binary black hole (BBH) problems, and this paper discussing an extension which couples GRMHD to dynamical spacetimes. In Section 2, we briefly discuss the GRMHD equations. Section 3 provides an overview of our numerical methods, with particular attention given to our atmosphere treatment and the implementation of a first-order flux correction (FOFC) scheme. We demonstrate the code’s accuracy and robustness in Section 4, including a convergence test with the first published BNS mergers performed on GPUs. Lastly, Section 5 describes the results of our performance and scaling tests, demonstrating AthenaK’s suitability for exascale machines. Lastly, we summarize our work in Section 6 and highlight some potential new NR applications for AthenaK.
@@ 3 Numerical Methods
In this section, we provide an overview of the numerical methods we use for the Valencia GRMHD solver. For an explanation of the refinement structure and methods shared with other modules in AthenaK, we refer the reader to Stone et al. (2024).
@@ 3 $L$ $5/3$ $1$ $10^{3}$ 00 00 00 $10$ $7$ $7$
3.4 Primitive Variable Recovery
4 Tests
In this section, we describe a number of tests to validate AthenaK.
4.1 Magnetized Shock Tubes
Figure 1: The rest-mass density $\rho$, pressure $P$, $x$-velocity $u^{x}=Wv^{x}$, and $y$-component of the magnetic field $B^{y}$ for the shock tube tests at $t=0.5$.
Our computational domain contains $1600$ cells. We use the LLF Riemann solver with PPM4 reconstruction. Because the fluid solver remains stable during this test and never approaches atmosphere, we do not enable the FOFC scheme. We evolve these systems to $t=0.5$ and compare them to their corresponding exact solutions (Giacomazzo & Rezzolla, 2006) in Figure 1. For brevity, we only show the $x$-component of velocity and the $y$-component of the magnetic field. The results are in good agreement with the exact solution.
@@ 4.2 Magnetized Cylindrical Blast Wave
For this test, we set $\rho_{\textrm{amb}}=10^{-4}$, $P_{\textrm{amb}}=3\times 10^{-5}$, $\rho_{\textrm{in}}=10^{-2}$, $P_{\textrm{in}}=1$, $r_{\textrm{in}}=0.8$, and $r_{\textrm{out}}=1.0$. We consider three values for the magnetic field: $B_{x}=0.01$, $B_{x}=0.1$, and $B_{x}=1.0$. The EOS is an ideal gas with $\Gamma=4/3$. As noted in Beckwith & Stone (2011), this test is particularly difficult for large $B_{x}$ because the combination of relatively high Lorentz factors and high magnetizations heavily often causes the conserved-to-primitive inversion to return poor-quality solutions (if it finds them at all). Common solutions to this issue involve altering the test (e.g., increasing the ambient pressure as in Leismann et al. (2005); Del Zanna et al. (2007); Beckwith & Stone (2011); Deppe et al. (2023)) or employing schemes which violate energy conservation (Komissarov, 1999; Mignone & Bodo, 2006; Phillips & Komissarov, 2024; Komissarov & Phillips, 2024). Consequently, the original setup as in Komissarov (1999) tests both the robustness of our primitive recovery routine and the ability of the FOFC and FOFC+DMP schemes to add enough dissipation to limit unphysical behavior.
@@
We show the results for the $B_{x}=0.01$ case in Figure 3, the $B_{x}=0.1$ case in Figure 4, and $B_{x}=1.0$ in Figure 5. Both the FOFC and FOFC+DMP schemes are in good agreement for the weakly and moderately magnetized cases. The strongly magnetized case exhibits significant noise for both solutions, with FOFC+DMP offering somewhat better behavior near the center. Nevertheless, the overall solution agrees with Komissarov (1999), and our results are consistent with what was found for the same test using the HARM-like solver (Stone et al., 2024).
@@ 4.3 Magnetic Loop Advection
Figure 6: (Top) The $B^{x}$ component of the magnetic field during the loop advection test. (Bottom) The violation of the divergence-free constraint relative to the maximum magnetic field strength. In both plots, the left half shows the initial data, and the right half shows the evolved data after a single period ($t=2.4$).
This test probes the correctness and the diffusiveness of the constrained-transport scheme. As shown in Figure 6, AthenaK advects the loop at the correct speed. Some numerical diffusion is apparent, but the shape is correct with no obvious artifacts. Maximum relative errors in the divergence-free condition are around $7\times 10^{-14}$ and primarily concentrated around the edge of the loop. After a single period, the maximum error has amplified to approximately $3\times 10^{-12}$, which is consistent with growth from floating-point errors.
@@ 4.4 Oscillating Neutron Star in the Cowling Approximation
The computational domain spans $[0~{}\textrm{M}_{\odot},102.4~{}\textrm{M}_{\odot}]$ in all dimensions with vacuum boundary conditions and a reflection symmetry along the $x$, $y$, and $z$ axes. The base refinement level has $128$ cells in each direction, and we use static mesh refinement (SMR) to refine all mesh blocks inside $51.2~{}\textrm{M}_{\odot}$ and then again at $25.6~{}\textrm{M}_{\odot}$ to give the innermost refinement region a resolution of $0.2~{}\textrm{M}_{\odot}$. We set the Courant-Friedrich-Lewy (CFL) factor to $\textrm{CFL}=0.4$.
@@ 4.6 Magnetized Accretion Disk
Figure 9: Diagnostic fluxes for the $64^{3}$, $96^{3}$, and $128^{3}$ cases plotted against some of the $96^{3}$, $128^{3}$, and $192^{3}$ EHT data (respectively). From top to bottom in each set of plots, we show the rates for mass accretion, normalized energy accretion, normalized angular momentum accretion, and normalized magnetic flux.
@@
We show these diagnostic quantities in Figure 9 for each run. For comparison, the plots also contain a subset of the data from Porth et al. (2019) (hereafter referred to as the EHT data), namely Athena++, BHAC, Cosmos++ (Anninos et al., 2005; Fragile et al., 2012, 2014), ECHO, HARM-Noble (usually called HARM3d) and iharm3D (Gammie et al., 2003; Noble et al., 2006, 2009), IllinoisGRMHD, and KORAL (Sadowski et al., 2013, 2014). Somewhat confusingly, we compare the $64^{3}$ AthenaK runs to the $96^{3}$ EHT data, the $96^{3}$ runs to the $128^{3}$ EHT data, and the $128^{3}$ runs to the $192^{3}$ EHT data. This is because the AthenaK grids do not have a direct correspondence with any setup in the EHT data; all AthenaK setups are Cartesian, while most of the EHT data setups are on logarithmic spherical grids. Nevertheless, we see good agreement between AthenaK and the other codes present at all resolutions, particularly for the angular momentum flux and the mass accretion rates.
@@ 4.7 Binary Neutron Star Merger
We use the SR and HR resolution runs to establish convergence of the waveform phase. To eliminate phase offsets caused by non-converging noise prior to the arrival of the so-called “junk” radiation at the extraction radius, we align the waveforms using the procedure in Boyle et al. (2008) between $t-r_{\ast}=0$ and the merger time of the LR waveform. We show the results in the bottom plot in Figure 15. Throughout the inspiral and merger, we maintain approximate second-order convergence as expected by our numerical methods. The results do not show clean convergence in the post-merger phase even before collapse. However, we note that most NR codes struggle to maintain convergence in the post-merger phase, especially with second-order schemes (Bernuzzi et al., 2012b; Radice et al., 2015; Most et al., 2019; Zappa et al., 2023). Though going to higher resolution may improve convergence, achieving second-order convergence is unlikely due to the presence of fluid turbulence, which, without including an explicit viscous term, will cascade to smaller length scales as resolution is increased.
@@
We next consider the reduced binding energy, $E_{b}=(M_{\mathrm{ADM}}-E_{\mathrm{GW}}-M)/(M\nu)$, as a function of the reduced angular momentum, $j=(J_{\mathrm{ADM}}-J_{\mathrm{GW}})/(M^{2}\nu)$, where for our initial data $M_{\mathrm{ADM}}=2.9984~{}\textrm{M}_{\odot}$, $J_{\mathrm{ADM}}=8.83542~{}\textrm{M}_{\odot}^{2}$ in natural units, and the symmetric mass ratio is $\nu=q/(1+q)^{2}=1/4$. This is a gauge-invariant measure which shows the evolution of the binary as it loses both energy and angular momentum due to gravitational radiation (Damour et al., 2012; Bernuzzi et al., 2012a). The bottom plot in Figure 16 shows that our runs are all in excellent agreement prior to merger. The lines do diverge somewhat in the post-merger phase, but the differences in the AthenaK runs decrease with resolution, indicating some measure of convergence.
@@ 5.2 Strong Scaling
We show the strong-scaling test results in Figure 19. Relative to a single node (${\sim}236^{3}$ zones per GPU), AthenaK on Frontier maintains $\geq 80\%$ efficiency down to ${\sim}96^{3}$ zones per GPU for the TOV test. The $64^{3}$ block test is roughly $45\%$ faster than the $32^{3}$ test, suggesting that $32^{3}$ mesh blocks do not saturate the GPUs properly. Despite the performance difference, scaling results are comparable between the two setups at smaller node counts. However, there is some evidence suggesting that large mesh blocks scale more poorly at large node counts. Though the limited tests shown here are hardly conclusive, this seems reasonable; larger mesh blocks improve GPU vectorization and reduce the total number of communication calls, but the cost of launching a kernel does not substantially change, meaning that kernel overhead will become a bottleneck sooner than in the case of small mesh blocks.
@@ 5.3 Discussion on Performance
{ruledtabular}
Table 2: Summary of AthenaK performance for a freely evolving TOV on different architectures. Speedup is measured relative to a single 64-core CPU socket on an AMD EPYC 7763. For reference, GR-Athena++ results on this system are also included.
Architecture $10^{6}$ zone-cycles/s Speedup
EPYC 7763 $0.760$ $1$
EPYC 7763 (GR-Athena++) $0.964$ $1.27$
A100 $10.7$ $14.1$
MI250X $4.84$ $6.37$
Ryzen 9 5900X $0.125$ $0.164$
RTX 3070 (laptop) $1.08$ $1.42$
In addition to the scaling results above, we performed a small number of tests comparing CPU and GPU performance on NERSC Perlmutter, which consists of CPU nodes (each with two AMD EPYC 7763 CPUs) and GPU nodes (each with a single CPU and four Nvidia A100 GPUs). Our setup is similar to Sec. 4.5 but with all output disabled and limited to the first 20 time steps. We also reduce the resolution to $128^{3}$ with mesh blocks of size $32^{3}$ to ensure they fit on a single 40 GB GPU. Running with a single CPU (64 cores/128 threads) achieves $7.60\times 10^{5}$ zone-cycles/s, and a single Nvidia A100 GPU manages to achieve $1.07\times 10^{7}$ zone-cycles/s, a factor of ${\sim}14$ improvement. The same setup on OLCF Frontier using a single MI250X compute die achieves $4.84~{}\times 10^{6}$ zone-cycles/s. To provide a baseline comparison, we also perform this test with GR-Athena++ using as similar a setup as possible. For a cell-centered spacetime, GR-Athena++ achieves $9.64\times 10^{5}$ zone-cycles/s, which is somewhat faster but comparable to AthenaK’s CPU performance.
We also ran a version of this test on a consumer-grade laptop with an Nvidia RTX 3070 laptop GPU. Due to the reduced amount of memory, we reduced the resolution by half, switched to $16^{3}$ mesh blocks, and imposed a reflection symmetry across the z-axis, but the physics options are the same. The laptop GPU achieves $1.08\times 10^{6}$ zone-cycles/s, which is $42\%$ faster than a single CPU on Perlmutter. The same calculation using the laptop’s CPU (an AMD Ryzen 9 5900X) with 16 OpenMP threads only achieved $1.25\times 10^{5}$ zone-cycles/s. All of these results are summarized in Table 2. We also found it possible to run a very coarse (${\sim}900~{}\mathrm{m}$) BNS simulation through merger on this same device in a matter of a few hours. Though research-grade calculations still require larger machines, the order-of-magnitude speedup provided by GPUs makes it possible to accelerate testing and debugging considerably.
@@ 6 Conclusion
The authors thank the Penn State Numerical Relativity Group for helpful discussions and important feedback. This research makes use of data from the EHT Code Comparison Project, and JF thanks Oliver Porth for assistance with accessing the data. Part of this work was completed at the TACC Open Hackathon, part of the Open Hackathons program. The authors would like to acknowledge OpenACC-Standard.org for their support. We are particularly grateful to Matthew Cawood, Victor Eijkhout, and Forrest Glines for helpful discussions while optimizing AthenaK. SB and BD acknowledge support by the EU Horizon under ERC Consolidator Grant, no. InspiReM-101043372. This research was supported by funding from the U.S. Department of Energy, Office of Science, Division of Nuclear Physics under Award Number(s) DE-SC0021177 and DE-SC0024388, by NASA under award No. 80NSSC21K1720, and by the National Science Foundation under Grants No. PHY-2011725, PHY-2020275, PHY-2116686, and AST-2108467. Part of this work was completed during a sabbatical visit to the Institute for Gravitation and the Cosmos at the Pennsylvania State University by SB. Simulations were performed on OLCF’s Frontier, NERSC’s Perlmutter, and on the Pennsylvania State University’s Institute for Computational and Data Sciences’s Roar Collab supercomputer. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.